Form a polynomial f(x) with real coefficients having the given degree and zeros.
Degree 5; Zeros:2; -i; -6+i

Answer by swincher4391(1107)

FACT: All imaginary roots come in conjugate pairs.
So we have 2,-i,i, -6+i, -6-i.
Then (x-2)[(x+i)(x-i)][(x-(-6+i))(x-(-6-i))] = (x-2)(x^2+1)(x^2+12x+37) = x^5+10x^4+14x^3-64x^2+13x-74